# TempFlow-GRPO: When Timing Matters for GRPO in Flow Models

### v1

@@ 4.2 Policy Gradient-Based Theoretical Justification
To provide a deeper understanding of our approach, we now examine it from the policy gradient perspective. Notice that in the $\sum_{k=0}^{T-1}$, k is the timestep, and in the equation, k is the value of the timestep. Simplifying Eq. 5, we obtain $\bm{x}_{k-1}\sim\mathcal{N}(\mu_{\theta}(\bm{x}_{k},k),\sigma_{k}^{2}\Delta k\bm{I})$, where:
@@
Starting from the policy gradient formulation in Eq. 7, we have:
@@
Eq. 26 indicates that the norm of $\mathbb{E}_{\bm{\epsilon}}[\bm{\epsilon}\hat{A}_{k}]$ is invariant among the timesteps. So in Eq. 22 and Eq. 23, the scale terms that modulate the contribution of each timestep’s model gradient $\nabla_{\theta}\bm{v}_{\theta}(\bm{x}_{k},k)$ to the overall reward gradient simplify to distinct functions, which we denote as $\sqrt{\frac{\Delta k(1-k)}{k}}$ and $\Delta k$. We visualize these scale terms under different flow shift, as depicted in the right of Fig. 4. As our analysis indicates, the setting of “w/o reweighting” exhibits a highly imbalanced contribution from each timestep’s model gradient to the final reward gradient. This non-uniformity systematically causes the contribution from the early denoising step where the model performs broad, structural exploration to be significantly smaller than the contribution from the late step, which focuses on fine-grained refinement. By employing our proposed noise-aware policy reweighting, this issue is substantially mitigated, as the scale term simplifies to being directly proportional to the step size, $\Delta k$. Furthermore, when the flow shift is set as 1, our method achieves a perfect equilibrium: it ensures that the gradient contribution from every single timestep is precisely equal, thereby completely balancing this effect across the entire generation trajectory.
5 Experiment
Following Flow-GRPO, we validate our approach on Compositional Image Generation in Geneval Ghosh et al. (2023) and Human Preference Alignment in PickScore Kirstain et al. (2023). To ensure a fair comparison, we normalized the weights applied to the policy loss to have a mean of 1 at all timesteps. We use num groups as 48 and group size as 24. The beta of kl loss, is set to 0.001 for PickScore and 0.004 in Geneval. Unless otherwise specified, the image size is set to 512, consistent with the configuration used for Flow-GRPO. It is worth noting that we further improve upon the original Flow-GRPO method by proposing a variant, Flow-GRPO (FIXED), which demonstrates superior performance. In our PickScore experiments, we adopt this improved version to ensure a fair and strong baseline for comparison.
5.1 Main Results
Model Step Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding
Diffusion Models
LDM Rombach et al. (2022) - 0.37 0.92 0.29 0.23 0.70 0.02 0.05
SD1.5 Rombach et al. (2022) - 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD2.1 Rombach et al. (2022) - 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL Podell et al. (2023) - 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 Ramesh et al. (2022) - 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 Betker et al. (2023) - 0.67 0.96 0.87 0.47 0.83 0.43 0.45
Autoregressive Models
Show-o Xie et al. (2024b) - 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Emu3-Gen Wang et al. (2024) - 0.54 0.98 0.71 0.34 0.81 0.17 0.21
JanusFlow Ma et al. (2025) - 0.63 0.97 0.59 0.45 0.83 0.53 0.42
Janus-Pro-7B Chen et al. (2025) - 0.80 0.99 0.89 0.59 0.90 0.79 0.66
GPT-4o Hurst et al. (2024) - 0.84 0.99 0.92 0.85 0.92 0.75 0.61
Flow Matching Models
FLUX.1 Dev Black et al. (2025) - 0.66 0.98 0.81 0.74 0.79 0.22 0.45
SD3.5-L Esser et al. (2024) - 0.71 0.98 0.89 0.73 0.83 0.34 0.47
SANA-1.5 4.8B Xie et al. (2025) - 0.81 0.99 0.93 0.86 0.84 0.59 0.65
SD3.5-M Esser et al. (2024) - 0.63 0.98 0.78 0.50 0.81 0.24 0.52
GRPO based Methods
SD3.5-M+Flow-GRPO Liu et al. (2025) 4400 0.90 0.99 0.97 0.90 0.88 0.85 0.80
SD3.5-M+TempFlow-GRPO 4400 0.97 1.00 0.99 0.96 0.97 0.98 0.90
Compositional Image Generation. We evaluate the compositional image generation capability of TempFlow-GRPO on the Geneval benchmark with its corresponding reward model. The experimental results are summarized in Tab. 1. As shown, our approach significantly improves the performance of the base model, increasing the overall score from 0.63 to 0.97. Furthermore, among GRPO-based methods, our method substantially outperforms Flow-GRPO: it achieves a performance of 0.97 within only 4,400 steps, whereas Flow-GRPO reaches only 0.90 under the same conditions. Additionally, as illustrated in Fig. 2, our method requires only about 2,000 steps to achieve a score of 0.95, while Flow-GRPO needs approximately 5,600 steps to reach the same level. Overall, these results demonstrate that TempFlow-GRPO not only accelerates convergence but also achieves superior final performance compared to existing approaches.
Human Preference Alignment. To further validate the generalizability of our approach, we conducted experiments on the PickScore benchmark, using PickScore as the reward model. As shown in the left of Fig. 2, our method, TempFlow-GRPO, achieves the highest performance, surpassing the original Flow-GRPO by approximately 1.3% and outperforming the improved baseline Flow-GRPO (Fixed) by about 0.6 %. Notably, our method requires only 100–200 training steps to match the performance of Flow-GRPO, and just 300–400 steps to reach the level of Flow-GRPO (Fixed). These results on PickScore further demonstrate the general applicability of our method as a unified flow-based RL algorithm across different reward models.
5.2 Analysis
Figure 5: Ablation Studies on trajectory branch and noise-aware policy reweighting. Left: PickScore Benchmark. Right: Geneval Benchmark.
Ablation of TempFlow-GRPO. We conducted ablation studies to explore the effectiveness of our proposed components: the trajectory branch and noise-aware policy weighting. These ablations were performed on both the Geneval and PickScore benchmarks. As shown in the Fig. 5, on the PickScore benchmark, introducing the trajectory branch further improves the performance of Flow-GRPO (Fixed), and applying noise-aware reweighting on top of this yields the highest pickscore results. On the Geneval benchmark, the benefit of the noise-aware strategy is even more significant: compared to Flow-GRPO, noise-aware policy reweighting boosts performance from 0.85 to 0.94, a 9% improvement while the trajectory branch also brings about a substantial gain of approximately 5%. These ablation results clearly demonstrate the effectiveness of our proposed methods.
Figure 6: Comparsion on PickScore benchmark in 1024 resolution.
Result of 1024 resolution. We further explored the effectiveness of our approach across different image resolutions, using PickScore as the reward model. As illustrated in the Fig. 6, after 450 training steps, TempFlow-GRPO achieves a 0.6% improvement in PickScore. Notably, our method requires only about 180 steps to reach the performance that Flow-GRPO (Fixed) attains after 450 steps. This further demonstrates the efficiency and effectiveness of our proposed method across varying resolutions.
Qualitative Result. We also conducted qualitative analyses on the SD3.5-M, Flow-GRPO (Fixed), and TempFlow-GRPO. As shown in the visualization results in Fig 7, compared to Flow-GRPO (Fixed), TempFlow-GRPO produces images with noticeably finer details and fewer visual artifacts or mistakes. In particular, our approach demonstrates superior capability in preserving complex structures and realistic textures. These qualitative improvements further highlighting the advantages of our method in generating high-quality, visually appealing images.
@@ 5.2 Analysis
Figure 7: Qualitative comparison between SD3.5-M, Flow-GRPO and TempFlow-GRPO with PickScore as reward. We use the red box to indicate the error region.
6 Limitation and Future Work
Although our method achieves significant improvements in both performance and image quality, the current experiments are based on a single reward model. In future work, we plan to focus on incorporating multi rewards from more powerful models, aiming to enhance performance across multiple dimensions. Additionally, we intend to design a comprehensive pipeline that can systematically improve various aspects of generative quality.
7 Conclusion
We presented TempFlow-GRPO, a temporally-aware reinforcement learning framework that addresses fundamental limitations in existing flow-based GRPO methods. Our key insight is that the uniform treatment of all timesteps creates a misalignment between optimization effort and actual impact on generation quality. Through trajectory branching, we enable precise credit assignment to intermediate actions without requiring specialized process reward models. Through noise-aware weighting, we ensure that optimization intensity matches each timestep’s exploration potential, preventing both under-exploration of critical early decisions and over-optimization of minor refinements. Our extensive experiments demonstrate that TempFlow-GRPO achieves state-of-the-art performance on human preference alignment and compositional generation benchmarks, surpassing both traditional GRPO methods and recent flow-based approaches.

### v2

@@ A.1 Policy Gradient-Based Theoretical Framework
To provide a deeper understanding of our approach, we now examine it from the policy gradient perspective. Note that in the summation $\sum_{k=0}^{T-1}$, the index $k$ denotes the timestep, while in subsequent equations, $k$ represents the timestep value. Simplifying Equation 5, we obtain $\bm{x}_{k-1}\sim\mathcal{N}(\mu_{\theta}(\bm{x}_{k},k),\sigma_{k}^{2}\Delta k\bm{I})$, where:
@@
Starting from the policy gradient formulation in Equation 8, we have:
@@
Equation 28 indicates that the norm of $\mathbb{E}_{\bm{\epsilon}}[\bm{\epsilon}\hat{A}_{k}]$ is invariant among the timesteps.
A.2 Experimental Setting Details
Dataset. We evaluate Compositional Image Generation on Geneval and Human Preference Alignment on PickScore. Meanwhile, we also test our method on HPDv2 Wu et al. (2023).
Reward Model. In TempFlow-GRPO, we introduce four reward models, including clip-based PickScore, HPSv2 Wu et al. (2023), VLM-based HPSv3 Ma et al. (2025b) and framework Geneval.
Compositional Image Generation. Following Flow-GRPO, we evaluate our method on PickScore prompts using the PickScore reward function. PickScore is a CLIP-based scoring function that demonstrates superhuman performance in predicting human preferences for generated images. To ensure consistency with Flow-GRPO, we set the KL divergence regularization coefficient $\beta=0.001$.
Human Preference Alignment. We adopt the GenEval framework following Flow-GRPO’s experimental protocol. The training prompts are sourced from the Flow-GRPO dataset. GenEval provides an object-focused evaluation framework that assesses compositional image properties, including object co-occurrence, spatial positioning, object count, and color attributes. For fair comparison, we maintain the KL divergence regularization coefficient at $\beta=0.004$.
HPSv3 Evaluation on FLUX.1-dev. For experiments on the FLUX.1-dev model, we configure the noise level at 0.9 for both Flow-GRPO and our TempFlow-GRPO method. The KL divergence weight is set to 0.004. During training, we employ 10 sampling steps, while evaluation uses 50 steps to ensure high-quality generation. The batch configuration consists of a group size of 24 (arranged as $4\times 6$) with 48 groups in total.
Comparison with DanceGRPO. DanceGRPO and Flow-GRPO represent concurrent developments in this research area. The SDE sampling methodology employed in DanceGRPO corresponds to Equation 12 in the Flow-GRPO formulation. However, the theoretical derivation in DanceGRPO terminates at its Equation 7 without developing an analogous final formulation. To comprehensively validate our method’s effectiveness, we conduct a direct comparison with DanceGRPO. For equitable comparison, we train TempFlow-GRPO for 300 iterations on the HPDv2 dataset using the HPSv2 reward function.
A.3 Experiments on FLUX.1-dev
[…]
Figure 11: The performance and KL loss (smoothed) of HPSv3 on HPDv2 dataset with FLUX.1-dev as base model.
A.4 Comparsion with DanceGRPO
[…]
A.5 Experiments on SD3.5-M with $\times1024$ resolution
[…]
Figure 12: (Left) Performance of TempFlow-GRPO on HPDv2 with HPSv2 as reward compared DanceGRPO. (Right) Comparsion on PickScore benchmark (PickScore reward, FLUX.1-dev 1024).
A.6 Training Hours
(a) Training time comparison on GenEval benchmark (Geneval reward, SD3.5-M 512).
(b) Training time comparison on PickScore prompts (PickScore reward, SD3.5-M 1024).
(c) Training time comparison on PickScore prompts (HPSv3 as reward, FLUX.1-dev 1024).
(d) Training time comparison on HPDv2 prompts (HPSv3 reward, FLUX.1-dev 1024).
Figure 14: Comprehensive comparison of training time across multiple benchmarks and resolutions. TempFlow-GRPO consistently achieves superior performance with reduced training time compared to Flow-GRPO baseline.
(a) Group strategy analysis on PickScore prompts (PickScore as reward, SD3.5-M 1024).
(b) Group strategy analysis on PickScore prompts (HPSv3 as reward, FLUX.1-dev 1024).
(c) Group strategy analysis on HPDv2 prompts (HPSv3 as reward, FLUX.1-dev 1024).
Figure 15: Comprehensive analysis of group strategy impact across different models and benchmarks, demonstrating consistent performance improvements with our proposed method and seed group approach.
A discussion of the computational cost of our method is warranted. Due to the $K$-branch exploration at each timestep, our sampling process incurs higher computational overhead compared to Flow-GRPO. Specifically, for $K=10$, the average number of branches is approximately $(9+\ldots+1)/10=4.5$ times that of Flow-GRPO. However, the training time per iteration remains identical to Flow-GRPO. Despite this increased per-sample computational cost, our method demonstrates substantially superior overall training efficiency. As illustrated in Figure 4 (bottom left, main paper) and Figure 14, where we plot performance against wall-clock training time, our method exhibits faster convergence across all evaluated benchmarks. This efficiency gain becomes particularly pronounced at higher resolutions, where our approach achieves Flow-GRPO’s final performance using only 33% to 50% of the total training time.
Furthermore, our approach not only converges faster but also achieves superior final performance. The PickScore experiments presented in the main paper reveal that even after extended training of 1200 steps, our method continues to outperform Flow-GRPO, providing compelling evidence that our approach possesses a higher performance ceiling.
A.7 Additional Details on Group Strategy
In Section 4.3 of the main paper, we analyze the impact of the group strategy on performance. Figure 8 and Figure 15 demonstrate that TempFlow-GRPO achieves substantial improvements even when employing an identical group strategy to the baseline, indicating the inherent advantages of our proposed core modules. To systematically disentangle the contributions of the group strategy from our two fundamental components—trajectory branching and noise reweighting—we conduct controlled ablation studies. In these experiments, we maintain a constant group strategy configuration to ensure fair comparison. The results unequivocally demonstrate that our core modules independently yield significant performance gains. These components not only enable the model to match Flow-GRPO’s multi-step performance in substantially fewer iterations but also elevate the final performance ceiling, validating the effectiveness of our theoretical framework.
A.8 KL Divergence Analysis
(a) KL Loss on PickScore (SD3.5-M 512).
(b) KL Loss on PickScore (SD3.5-M 1024).
Figure 16: Comprehensive analysis of KL divergence dynamics during training across different resolutions. TempFlow-GRPO consistently maintains lower KL divergence values compared to Flow-GRPO, indicating better preservation of the original model distribution while achieving superior preference alignment.
To further substantiate the training stability of our method, we provide a comprehensive analysis of the KL divergence dynamics during optimization. The results, presented in Figure 16, clearly demonstrate that our method consistently maintains significantly lower KL divergence values throughout the training process compared to Flow-GRPO. This reduced divergence indicates better preservation of the original model distribution while still achieving effective preference alignment.
Furthermore, we analyze the impact of different grouping strategies on KL divergence. While adopting the “seed group” strategy results in a marginal increase in KL divergence compared to our default configuration, the values remain substantially below those of the Flow-GRPO baseline across all training iterations. This observation provides strong evidence that our core modules—trajectory branching and noise reweighting—inherently enhance training stability and robustly maintain distributional proximity, regardless of the specific grouping strategy employed. The consistent stability across different configurations underscores the fundamental advantages of our theoretical framework in balancing exploration and exploitation during the optimization process.
A.9 Multi-Reward Optimization
Table 2: Performance comparison for multi-reward optimization on PickScore prompts, demonstrating simultaneous improvement across multiple objectives.
    HPSv3 $\uparrow$     PickScore $\uparrow$
    FLUX.1-dev     9.42     22.34
    Flow-GRPO (Prompt)     10.41 (+0.99)     22.44 (+0.1)
    TempFlow-GRPO     12.23 (+1.81)     22.52 (+0.18)
We evaluate the performance of our method in multi-reward optimization scenarios. For this experiment, we employ FLUX.1-Dev as the base model and train at a high resolution of $\times1024$. We simultaneously optimize for two reward functions—HPSv3 and PickScore—with a weight ratio of 1:0.26. After 120 training steps, the results presented in Table 2 demonstrate our method’s robust capability for multi-objective optimization. Compared to Flow-GRPO, our approach achieves a substantial improvement of 0.82 on the HPSv3 metric while simultaneously maintaining consistent performance gains on PickScore. These results validate that our temporal reweighting mechanism effectively balances multiple objectives without sacrificing performance on individual metrics, highlighting the versatility of our approach in complex preference alignment scenarios where multiple criteria must be optimized concurrently.
A.10 Visualization
We present comprehensive qualitative comparisons to demonstrate the superior visual quality achieved by our TempFlow-GRPO method. The visualization results consistently show that our approach generates images with enhanced fidelity, better adherence to complex prompts, and fewer visual artifacts compared to baseline methods.
Prompts in Figure 1. The prompts in Figure 1 are as follows:
1. 16-year-old teenager wearing a white bear-ear hat with a smirk on their face. 2. photo of well done salmon dinner, 8K, Global Illumination, Ray Tracing Reflections 3. A lemon with a McDonald’s hat. 4. An image of an emo with dark brown hair in a messy pixie cut, large entirely-black eyes, wearing black clothing. 5. The image is a mixed media collage with broken glass and torn paper elements, featuring intricate oil details and a canvas texture, in a contemporary art style. 6. An epic deep space photograph. Gigantic, monolithic letters forming the word ’TempFlow-’ hang silently in the void, their surfaces like ancient, cracked obsidian reflecting distant starfields. Far below, the letters ’GRPO’ are formed by a vast, tranquil nebula glowing with soft, ethereal light, like a cosmic ocean. The sense of scale is immense and humbling. Perfect, case-sensitive lettering. Moody, atmospheric, photorealistic, cinematic wide-angle shot, 4K UHD. 7. Kiwi fruit, mint leaves, ice cubes, background yellow, splashing water, soft box, back light, creative food photography, Art by Alberto Seveso, 8. Claymation of Futurama characters. 9. A group of four friends commemorating a ski trip in the snow. 10. an empty bench next to a busy street. photo of well done salmon dinner, 8K, Global Illumination, Ray Tracing Reflections 11. a 12 year old girl and her pet raccoon
@@ A.10 Visualization
Figure 17: Qualitative comparison between FLUX.1-dev, Flow-GRPO (Prompt) and TempFlow-GRPO with HPSv3 rewards on HPDv2 prompts.
Figure 18: Qualitative comparison between FLUX.1-dev, Flow-GRPO (Prompt) and TempFlow-GRPO with multi rewards on PickScore prompts.
Figure 19: Qualitative comparison between FLUX.1-dev, Flow-GRPO (Prompt) and TempFlow-GRPO with multi rewards on PickScore prompts.
Figure 20: Qualitative comparison between FLUX.1-dev, Flow-GRPO (Prompt) and TempFlow-GRPO with multi rewards on PickScore prompts.
Figure 21: Qualitative comparison between FLUX.1-dev, Flow-GRPO (Prompt) and TempFlow-GRPO with multi rewards on PickScore prompts.